• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 12:06:31 on September19,2005

    Entry number 155197

    keyword=Q2 measurement

    Took Spectrometer DAQ runs to measure Q2 off of the Helium target.

    The Left-spectrometer IOC (iocha10) went down just as the runs were beginning. Since the magnet settings stayed the same and the beam was going away at 11am, we just forged on and took the data. It looks good.

    10:10 started switching over DAQ for Q2 measurement, following the checklist.

    Rates with 1uA, and 0.5uA (in Hz)

    1uA 0.5uA 0.2uA
    T1 823k 440k 182k
    T2 96k 45k 17k
    T3 780k 410k 169k
    T4 126k 62k 24k
    T5 1 0 1
    T6 0 0 0
    T7 880k 475k 254k
    T8 950k 480k 199k


    Runs were:

    2686 short (25k) events at 0.2uA.
    2687 Q2 run at 0.2uA, 1M events
    2688 short Q2 run at 0.2uA, 230k events.
    2689 Q2 run at 0.5uA, 1M events.
    2690 Q2 run at 0.5uA, 1M events.
    2691 very short run with double prescale factors of 2690. Beam taken away at the end. 100k events.


    11:05 Beam taken away for injector work. Switched back the HAPPEX mode.


    Quick analyses of runs 2687 (Figure 2), 2689, and 2690 (Figure 1) give the same general distributions for different beam currents and focal plane rates. Note that at 0.2uA, the position of the beam could not be measured by the stripline BPMs.



    FIGURE 1

    FIGURE 2